If it takes 17 minutes to read 102 words, how many minutes would it take you to read 222 words?
Answer by TimothyLamb(4379)

17/102 = x/222
x = 222*17/102
x = 37 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php